# MixGRPO: Unlocking Flow-based GRPO Efficiency with Mixed ODE-SDE

### v1

@@ 3.2 Sliding Window as Optimization Scheduler
In fact, the interval $S$ can be non-fixed during the training process. In this section, we will introduce the sliding window to describe the movement of the interval $S$, which leads to a significant improvement in the quality of the generated images. Along the discrete denosing timesteps $\{0,1,\ldots,T-1\}$, MixGRPO defines a sliding window $W(l)$ and optimization is only employed at the timesteps within this window.
$W(l)=\{t_{l},t_{l+1},\dots,t_{l+w-1}\},\quad l\leq T-w$ (10)
@@
The left boundary $l$ of the sliding window moves as the training progresses. In the experiments, we found that the window size $w$, shift interval $\tau$, and window stride $s$ are all crucial hyperparameters. Through ablation studies (See Experiment 4.4.1), we identified the optimal settings. When the total sampling steps $T=25$, the best performance is achieved with $w=4$, $\tau=25$ and $s=1$. The detailed sliding window strategy and the MixGRPO algorithm can be referenced in Algorithm 1
Restricting the use of SDE sampling within the sliding window not only ensures the diversity of the generated images but also allows the model to concentrate on optimizing the flow within that window. Moving along the denoising direction represents the stochasticity of the probability flow from strong to weak, as illustrated in Figure 2. This is essentially a greedy strategy and is similar to the RL that assigns discount factors to process rewards (Pitis, 2019; Amit et al., 2020; Hu et al., 2022), which gives greater importance to rewards derived from a larger search space in the earlier process. The experimental results of different movement strategies in Table 4 demonstrate the validity of this intuition. Not only does the progressive strategy outperform random selection, but we also found that even when the sliding window is not moved (frozen), meaning only the earlier timesteps are optimized, MixGRPO can still yield good results, particularly in terms of ImageReward (Xu et al., 2023) and UnifiedReward (Wang et al., 2025). Based on this intuition, we also proposed an exponential decay strategy as follows, allowing $\tau$ to decrease as the window moves, enabling the model to avoid excessive optimization in smaller search spaces.
@@
where $\tau_{0}$ is the initial shift interval, $k$ is the decay factor, and $\lambda_{thr}$ is the threshold that controls when the decay starts. The exponential function $\exp(x)$ calculates $e^{x}$, while the Rectified Linear Unit $\text{ReLU}(x)$ is defined as $\max(0,x)$. Table 4 shows that the exponential decay strategy can achieve better results in terms of Pick Score (Kirstain et al., 2023) and ImageReward (Xu et al., 2023), likely because the model focuses on the earlier stages of denoising, which can lead to more significant high-level changes.
@@ 3.3 Trade-off Between Overhead and Performance
Gao et al. (2024) has demonstrated the equivalence between the ODE sampling of flow matching models (FM) and DDIM, and Section 3.1 has also shown that diffusion probabilistic models (DPM) and FM share the same ODE form during the denoising process. Therefore, the higher-order ODE solvers e.g., DPM-Solver Series (Lu et al., 2022a; b; Zheng et al., 2023), UniPC (Zhao et al., 2023) designed for DPM sampling acceleration are also applicable to FM. We have reformulated DPM-Solver++ (Lu et al., 2022b) to apply it in the FM framework for ODE sampling acceleration and released detailed derivations in Appendix B.
By applying higher-order solvers, we achieve acceleration in the sampling of $\pi_{\theta_{\text{old}}}$ during GRPO training, which is essentially a balance between overhead and performance. Excessive acceleration leads to fewer timesteps, which inevitably results in a decline in image generation quality, thereby accumulating errors in the computation of rewards. We have found in practice that the 2nd-order DPM-Solver++ is sufficient to provide significant acceleration while ensuring that the generated images align well with human preferences. Ultimately, we adopted both progressive and frozen sliding window strategies, proposing MixGRPO-Flash and MixGRPO-Flash*. A detailed description of the algorithm can be found in Appendix C. These approaches achieve a greater degree of acceleration compared to MixGRPO, while also outperforming DanceGRPO in terms of performance.
@@
Method $\text{NFE}_{\pi_{\theta_{\text{old}}}}$ $\text{NFE}_{\pi_{\theta}}$ Iteration Time (s)$\downarrow$ Human Preference Alignment
@@ 4.1 Experiment Setup
Performance Evaluation  We used four reward models, namely HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), ImageReward (Xu et al., 2023) and Unified Reward Wang et al. (2025), both as reward guidance during training and as evaluation metrics. These metrics are all based on human preferences but emphasize different aspects. For example, ImageReward (Xu et al., 2023) highlights image-text alignment and fidelity, while Unified Reward (Wang et al., 2025) concentrates on semantics. DanceGRPO Xue et al. (2025) also demonstrates that using multiple reward models can achieve better results. To validate the robustness of MixGRPO, we also followed DanceGRPO and conducted additional comparisons using HPS-v2.1 as a single reward, and combining HPS-v2.1 (Wu et al., 2023) and CLIP Score (Radford et al., 2021) as multi-rewards.
@@ 4.2 Implementation Details
For training-time sampling, we first perform a shift $\tilde{t}=\frac{t}{1-(\tilde{s}-1)t}$ on the uniformly distributed $t=[0,\ldots,T-1]$ and then define $\sigma_{t}=\eta\sqrt{\frac{\tilde{t}}{1-\tilde{t}}}$. We set $\tilde{s}=3$ and $\eta=0.7$ as scale. We set $T=25$ as the total sampling steps. For GRPO, the model generates 12 images for each prompt and clips the advantage to the range [5,5][-5,5][ - 5 , 5 ]. It is important to note that we use gradient accumulation over 3 steps, which means that during a single training iteration, there are $\frac{12}{3}=4$ gradient updates. For the exponential decay strategy of the sliding window in Eq. ( 11), we empirically set $\tau_{0}=20$, $k=0.1$, and $\lambda_{thr}=13$. Furthermore, when multiple reward models are jointly trained with, each reward is assigned equal weight.
@@ 4.3 Main Experiments
Following DanceGRPO (Xue et al., 2025), we also trained and evaluated the model using a single reward model and two reward models, but on the HPDv2 dataset (Wu et al., 2023). The results (See Table 2) demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards, whether using a single or multiple reward models. The visualized results are displayed in Appendix D.
Table 2: Comparison results for in-domain and out-of-domain reward metrics. The results demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards, whether using a single or multiple reward models.
@@ 5 Conclusion
Although GRPO (Shao et al., 2024) has seen significant success in the language modality, it is still in the early stages of progress in vision (Tong et al., 2025; Xue et al., 2025; Liu et al., 2025). Existing flow-matching based GRPO faces challenges such as low sampling efficiency and slow training. To address these issues, we proposed MixGRPO, a novel training framework that combines SDE and ODE sampling. This hybrid approach allows for focused optimization on the SDE sampling flow component, reducing complexity while ensuring accurate reward computation. Inspired by the decay factor in reinforcement learning (Hu et al., 2022), we introduce a sliding window strategy for scheduling the optimized denoising steps. Experimental results confirm the effectiveness of our approach in both single-reward and multi-reward settings. Additionally, MixGRPO decouples the denoising stages for optimization and reward computation, enabling acceleration of the latter with high-order solvers. We further propose MixGRPO-Flash, which balances overhead and performance. We hope MixGRPO will inspire deeper research into post-training for image generation, contributing to the advancement of Artificial General Intelligence (AGI).
Appendix A Hybrid Inference for Solving Reward Hacking
@@
Appendix B DPM-Solver++ for Recitified Flow
The DPM-Solver++ algorithm (Lu et al., 2022b) is originally designed for the $\mathbf{x}_{0}\text{-}prediction$ diffusion model (Rombach et al., 2022), where the model outputs the denoised feature $\mathbf{x}_{0}$ based on the noisy feature $\mathbf{x}_{t}$, the time condition $t$, and the text condition $c$. According to the definition of Rectified Flow (RF) (Liu et al., 2022), there is the following transfer equation:
$\mathbf{x}_{t}=t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0}$ (12)
According to the theory of stochastic interpolation (Albergo et al., 2023), RF effectively approximates $\mathbf{x}_{1}-\mathbf{x}_{0}$ by modeling $\mathbf{v}_{t}$:
$\mathbf{v}_{t}=\mathbf{x}_{1}-\mathbf{x}_{0}$ (13)
@@ Appendix B DPM-Solver++ for Recitified Flow
$\mathbf{x}_{0}=\mathbf{x}_{t}-\mathbf{v}_{t}t$ (14)
@@
$\mathbf{x}_{\theta}=\mathbf{x}_{t}-\mathbf{v}_{\theta}t$ (15)
Taking the multistep second-order solver as an example (see Algorithm 2 in  (Lu et al., 2022b)), we derive the corrected $\mathbf{x}_{\theta}$ for the RF sampling process as $\mathbf{D}_{t}$:
$\mathbf{D}_{t}\leftarrow\left(1+\frac{h_{t}}{2h_{t-1}}\right)\left(\mathbf{x}_{t-1}-\mathbf{v}_{\theta}(\mathbf{x}_{t-1},t-1,c)\frac{t-1}{T}\right)$ (16)
$-\frac{h_{t}}{2h_{t-1}}\left(\mathbf{x}_{t-2}-\mathbf{v}_{\theta}(\mathbf{x}_{t-2},t-2,c)\frac{t-2}{T}\right)$
where $T$ is the total sampling steps and $h_{t}=\lambda_{t}-\lambda_{t-1}$. $\lambda_{t}$ is the log-signal-to-noise-ratio (log-SNR) and is defined in RF as $\lambda_{t}:=\log\left(\frac{1-t}{t}\right)$.
@@
$\mathbf{x}_{t}\leftarrow\frac{t}{t-1}\mathbf{x}_{t-1}-(1-t)\left(e^{-h_{t}}-1\right)\mathbf{D}_{t},\quad 1\leq t<T$ (17)
Appendix C MixGRPO-Flash Algorithm
MixGRPO-Flash accelerates the ODE sampling that does not contribute to the calculation of the policy ratio after the sliding window by using DPM-Solver++ in the Eq. (17). We introduce a compression rate $\tilde{r}$ such that the ODE sampling after the window only requires $(T-l-w)\tilde{r}$ time steps. And the total time-steps is $\tilde{T}=l+w+(T-l-w)\tilde{r}$ The final algorithm is as follows:
Algorithm 2 MixGRPO-Flash Training Process
1:initial policy model $\pi_{\theta}$; reward models $\{R_{k}\}_{k=1}^{K}$; prompt dataset $\mathcal{C}$; total sampling steps $\tilde{T}$; number of samples per prompt $N$; ODE compression rate $\tilde{r}$

### v2

@@ 3.2 Sliding Window as Optimization Scheduler
In fact, the interval $S$ can be non-fixed during the training process. In this section, we will introduce the sliding window to describe the movement of $S$, which leads to a significant improvement in the quality of the generated images. Along the denosing time-steps $\{0,1,\ldots,T-1\}$, MixGRPO defines a SDE sliding window $W(l)$ and optimization is only employed at the steps within $W(l)$.
$W(l)=\{t_{l},t_{l+1},\dots,t_{l+w-1}\},\quad l\leq T-w,$ (10)
@@
The left boundary $l$ of the sliding window moves as the training progresses. In the experiments, we found that the window size $w$, shift interval $\tau$, and window stride $s$ are all crucial hyperparameters. Through ablation studies (See Experiment 4.4.1), we identified the optimal settings. When the total sampling steps $T=25$, the best performance is achieved with $w=4$, $\tau=25$ and $s=1$. The detailed sliding window strategy and MixGRPO algorithm can be found in Algorithm 1.
Restricting the use of SDE sampling within the sliding window not only ensures the diversity of the generated images but also allows the model to concentrate on optimizing the flow within that window. Moving along the denoising direction represents the stochasticity of the probability flow from strong to weak, as illustrated in Figure 2. This is essentially a greedy strategy and is similar to the RL that assigns discount factors to process rewards (Pitis, 2019; Amit et al., 2020; Hu et al., 2022b), which gives greater importance to rewards derived from a larger search space in the earlier process.
The experimental results of different movement strategies in Table 5 demonstrate the validity of this intuition. Not only does the progressive strategy outperform random selection, but we also found that even when the sliding window is not moved (frozen), meaning only the earlier timesteps are optimized, MixGRPO can still yield good results, particularly in terms of ImageReward (Xu et al., 2023) and UnifiedReward (Wang et al., 2025). Based on this intuition, we also proposed an exponential decay strategy as follows, allowing $\tau$ to decrease as the window moves, enabling the model to avoid excessive optimization in smaller search spaces.
@@
where $\tau_{0}$ is the initial shift interval, $k$ is the decay factor, and $\lambda_{thr}$ is the threshold that controls when the decay starts. The exponential function $\exp(x)$ calculates $e^{x}$, while the Rectified Linear Unit $\text{ReLU}(x)$ is defined as $\max(0,x)$. Table 5 shows that the exponential decay strategy can achieve better results in terms of Pick Score (Kirstain et al., 2023) and ImageReward (Xu et al., 2023). This may be because the model focuses on the earlier stages of denoising, which can lead to more significant high-level changes, precisely what the human preference alignment reward model emphasizes.
@@ 3.3 Trade-off Between Overhead and Performance
Gao et al. (2024) has demonstrated the equivalence between the ODE sampling of flow matching models (FM) and DDIM, and Section 3.1 has also shown that diffusion probabilistic models (DPM) and FM share the same ODE form during the denoising process. Therefore, the higher-order ODE solvers e.g., DPM-Solver Series (Lu et al., 2022a; b; Zheng et al., 2023), UniPC (Zhao et al., 2023) designed for DPM sampling acceleration are also applicable to FM. We have reformulated DPM-Solver++ (Lu et al., 2022b) to apply it in the FM framework for ODE sampling acceleration and released detailed derivations in Appendix C.
By applying higher-order solvers, we achieve acceleration in the sampling of $\pi_{\theta_{\text{old}}}$ during GRPO training, which is essentially a balance between overhead and performance. Excessive acceleration leads to fewer timesteps, which inevitably results in a decline in image generation quality, thereby accumulating errors in the computation of rewards. We have found in practice that the 2nd-order DPM-Solver++ is sufficient to provide significant acceleration while ensuring that the generated images align well with human preferences in Table 9. Ultimately, we introduced DPM-Solver++ and adopted both progressive and frozen sliding window strategies, proposing MixGRPO-Flash and MixGRPO-Flash*. A detailed description of the algorithm can be found in Appendix E. These approaches achieve a greater degree of acceleration compared to MixGRPO, while also outperforming DanceGRPO in terms of human preference alignment performance.
@@
[…]
@@ 4.1 Experiment Setup
Performance Evaluation  We used four multiple reward models in conjunction for GRPO, namely HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), ImageReward (Xu et al., 2023) and Unified Reward Wang et al. (2025), both as reward guidance during training and as evaluation metrics. These metrics are all based on human preferences but emphasize different aspects. For example, ImageReward (Xu et al., 2023) highlights image-text alignment and fidelity, while Unified Reward (Wang et al., 2025) concentrates on semantics. DanceGRPO Xue et al. (2025) also demonstrates that using multiple reward models can achieve better results. To validate the robustness of MixGRPO, we also followed DanceGRPO and conducted additional comparisons using HPS-v2.1 as a single reward, and combining HPS-v2.1 (Wu et al., 2023) and CLIP Score (Radford et al., 2021) as multi-rewards.
@@ 4.2 Implementation Details
For training-time sampling, we first perform a shift $\tilde{t}=\frac{t}{1-(\tilde{s}-1)t}$ on the uniformly distributed $t_{i}=\frac{i}{T}$ where $i=[0,\ldots,T-1]$, and then define $\sigma_{t}=\eta\sqrt{\frac{\tilde{t}}{1-\tilde{t}}}$. We set $\tilde{s}=3$ and $\eta=0.7$ as scale. We set $T=25$ as the total sampling steps. For GRPO, the model generates 12 images for each prompt and clips the advantage to the range [5,5][-5,5]. It is important to note that we use gradient accumulation over 3 steps, which means that during a single training iteration, there are $\frac{12}{3}=4$ gradient updates. For the exponential decay strategy of the sliding window in Eq. ( 11), we empirically set $\tau_{0}=20$, $k=0.1$, and $\lambda_{thr}=13$. Furthermore, when multiple reward models are jointly trained with, each reward is assigned equal weight.
@@ 4.3 Main Experiments
Following DanceGRPO (Xue et al., 2025), we also trained and evaluated the model using a single reward model, e.g., HPSv2.1 and two reward models, e.g., HPSv2.1 and CLIP Score on the HPDv2 dataset (Wu et al., 2023). The results (See Table 3) demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards, whether using a single or multiple reward models. The visualized results are displayed in Figure 6 and Figure 7 of Appendix F.
To demonstrate the robustness of the MixGRPO, we followed Flow-GRPO (Liu et al., 2025a) and applied the Low-Rank Adaptation (LoRA) (Hu et al., 2022a) method on the Stable Diffusion 3.5 (SD3.5) (Esser et al., 2024). We used HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), and ImageReward (Xu et al., 2023) as multi-rewards, comparing Flow-DPO (offline/online) (Liu et al., 2025b), and Flow-GRPO (Liu et al., 2025a). The results in Table 3 showed that MixGRPO is well-suited for the LoRA, significantly reducing training overhead while outperforming both Flow-DPO and Flow-GRPO in terms of speed and performance. The visualization results can be found in Figure 8 of Appendix F.
Table 2: Comparison results demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards.
@@
Table 3: Comparison results demonstrate that MixGRPO outperforms Flow-DPO and Flow-GRPO in training efficiency and performance.
Model RL Method $\text{NFE}_{\pi_{\theta_{\text{old}}}}$ $\text{NFE}_{\pi_{\theta}}$ HPSv2.1 Pick Score ImageReward
SD3.5-M / / / 0.3066 0.2266 1.1630
SD3.5-M+DPO LoRA Offline DPO 40 40 0.3043 0.2220 1.4524
Online DPO 40 40 0.3132 0.2210 1.5001
SD3.5-M+GRPO LoRA Flow-GRPO 10 10 0.3312 0.2318 1.4572
MixGRPO 10 4 0.3416 0.2360 1.4854
@@ 5 Conclusion
Although GRPO (Shao et al., 2024) has seen significant success in the language modality, it is still in the early stages of progress in vision (Tong et al., 2025; Xue et al., 2025; Liu et al., 2025a). Existing flow-based GRPO faces challenges such as low sampling efficiency and slow training. To address these issues, we proposed MixGRPO, a novel training framework that combines SDE and ODE sampling. This hybrid approach allows for focused optimization on the SDE sampling flow component, reducing complexity while ensuring accurate reward computation. Inspired by the decay factor in reinforcement learning (Hu et al., 2022b), we introduce a sliding window strategy for scheduling the optimized denoising steps. Experimental results confirm the effectiveness of our approach in both single-reward and multi-reward settings. Additionally, MixGRPO decouples the denoising stages for optimization and reward computation, enabling acceleration of the latter with high-order solvers. We further propose MixGRPO-Flash, which balances overhead and performance. We hope MixGRPO will inspire deeper research into post-training for image generation, contributing to the advancement of Artificial General Intelligence (AGI).
6 Acknowledgement
The authors thank Yiming Cheng (M.Sc., Tsinghua University, 13306339695@163.com) for his valuable feedback on the theoretical derivations and mathematical expressions.
Appendix A Proof of convergence for Mixed ODE-SDE sampling
To prove that the mixed ODE-SDE sampling method in Eq. (4) has the same convergence as Eq. (2), which uses only ODE sampling, referencing Song et al. (2020b), we approach this from the perspective of distribution evolution, where the distribution at each time step, e.g., $\frac{\partial{q_{t}(\mathbf{x})}}{\partial{t}}$ must be the same. Let the interval for SDE be defined as $S=\left[t_{l},t_{r}\right)\in\left[0,1\right)$. Along the denoising direction, when the same initial Gaussian noise distribution $q_{0}(\mathbf{x}_{0})$ is given, the probability distribution evolution in the ODE interval preceding the SDE is completely identical. The key point is whether the distribution evolution of the SDE within the interval $S$ is completely equivalent to that of the ODE. If they are equivalent, then the ODE interval following the SDE will naturally be equivalent to using only ODE sampling. Next, we will provide a detailed proof for this key point.
Consider the SDE Eq. (3) in the interval $S$, which possesses the following form:
$\mathrm{d}\mathbf{x}=[f(\mathbf{x},t)-g^{2}(t)\nabla_{\mathbf{x}}\log{q_{t}(\mathbf{x})}]\mathrm{d}t+g(t)\mathrm{d}\mathbf{w},\quad t\in S.$ (12)
The marginal probability density $q_{t}(\mathbf{x}_{t})$ evolves according to Kolmogorov’s equation (Fokker-Planck equation) (Øksendal, 2003)
$\frac{\partial q_{t}(\mathbf{x})}{\partial t}=-\nabla_{\mathbf{x}}\cdot\big[\big(f(\mathbf{x},t)-g^{2}(t)\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\big)q_{t}(\mathbf{x})\big]+\frac{1}{2}g^{2}(t)\nabla^{2}_{\mathbf{x}}q_{t}(\mathbf{x})$ (13)
[…]
$\frac{\partial q_{t}(\mathbf{x})}{\partial t}$ $=-\nabla_{\mathbf{x}}\cdot\big[f(\mathbf{x},t)q_{t}(\mathbf{x})-g^{2}(t)\nabla_{\mathbf{x}}q_{t}(\mathbf{x})\big]+\frac{1}{2}g^{2}(t)\nabla^{2}_{\mathbf{x}}q_{t}(\mathbf{x})$
$=-\nabla_{\mathbf{x}}\cdot[f(\mathbf{x},t)q_{t}(\mathbf{x})-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}}q_{t}(\mathbf{x})]$
$=-\nabla_{\mathbf{x}}\cdot\big[\underbrace{\big(f(\mathbf{x},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\big)}_{f_{\text{ODE}}(\mathbf{x},t)}q_{t}(\mathbf{x})\big].$ (14)
The Eq. (A) is indeed the Fokker-Planck equation of the ODE Eq. (2). Therefore, within the interval $S$, the distribution evolution of SDE and ODE sampling is consistent.
Appendix B Hybrid Inference for Solving Reward Hacking
@@
Appendix C DPM-Solver++ for Recitified Flow
For clarity and to avoid ambiguity between continuous time and discrete steps, we adopt the following notation in this section. We denote the discrete time steps by an index $i\in\{0,1,\dots,T-1\}$, where $T$ is the total number of sampling steps. The continuous time corresponding to step $i$ is denoted by $t_{i}=\frac{i}{T}\in[0,1)$.
The DPM-Solver++ algorithm (Lu et al., 2022b) is originally designed for the $\mathbf{x}_{0}\text{-}prediction$ diffusion model (Rombach et al., 2022), where the model outputs the denoised feature $\mathbf{x}_{0}$ based on the noisy feature $\mathbf{x}_{t_{i}}$, the time condition $t_{i}$ and the text condition $c$. According to the definition of Rectified Flow (RF) (Liu et al., 2022), there is the following transfer equation:
$\mathbf{x}_{t_{i}}=t_{i}\mathbf{x}_{1}+(1-t_{i})\mathbf{x}_{0}.$ (15)
According to the theory of stochastic interpolation (Albergo et al., 2023), RF effectively approximates $\mathbf{x}_{1}-\mathbf{x}_{0}$ by modeling $\mathbf{v}_{t_{i}}$:
$\mathbf{v}_{t_{i}}=\mathbf{x}_{1}-\mathbf{x}_{0}.$ (16)
@@ Appendix C DPM-Solver++ for Recitified Flow
$\mathbf{x}_{0}=\mathbf{x}_{t_{i}}-\mathbf{v}_{t_{i}}t_{i}.$ (17)
@@
$\mathbf{x}_{\theta}(\mathbf{x}_{i},t_{i},c)=\mathbf{x}_{i}-\mathbf{v}_{\theta}(\mathbf{x}_{i},t_{i},c)\cdot t_{i}.$ (18)
Taking the multistep second-order DPMSolver++ as an example (see Algorithm 2 in  (Lu et al., 2022b)), we derive the corrected $\mathbf{x}_{\theta}$ for the RF sampling process as $\mathbf{D}_{i}$:
$\mathbf{D}_{i}\leftarrow\left(1+\frac{h_{i}}{2h_{i-1}}\right)\left(\mathbf{x}_{i-1}-\mathbf{v}_{\theta}(\mathbf{x}_{i-1},t_{i-1},c)\cdot t_{i-1}\right)$ (19)
$\quad-\frac{h_{i}}{2h_{i-1}}\left(\mathbf{x}_{i-2}-\mathbf{v}_{\theta}(\mathbf{x}_{i-2},t_{i-2},c)\cdot t_{i-2}\right),$
where $h_{i}=\lambda_{t_{i}}-\lambda_{t_{i-1}}$. The continuous time $t_{i}$ corresponds to the discrete step $i$ over a total of $T$ sampling steps. The term $\lambda_{t_{i}}$ is the log-signal-to-noise-ratio (log-SNR) and is defined in RF as:
$\lambda_{t_{i}}:=\log\left(\frac{1-t_{i}}{t_{i}}\right).$ (20)
@@
$\mathbf{x}_{i}\leftarrow\frac{t_{i}}{t_{i-1}}\mathbf{x}_{i-1}-(1-t_{i})\left(e^{-h_{i}}-1\right)\mathbf{D}_{i},\quad 1\leq i<T.$ (21)
Appendix D MixGRPO Algorithm
Taking the progressive movement strategy of the sliding window as an example, the MixGRPO Algorithm 1 is as follows. MixGRPO first sets the left boundary $l$ of the sliding window $W(l)$ at $t=0$ of the reverse process, which corresponds to the first sampling time-step. A GRPO iteration consists of two phases: sampling and optimization.
In the sampling phase, mixed ODE-SDE sampling is performed for the $N$ samples within the group, using SDE within the window and ODE outside. Then, the reward and advantage for the $N$ samples are calculated. For the multi-reward strategy, the advantage for each reward model is computed separately and then aggregated to obtain the final advantage. In the optimization phase, the GRPO loss (Shao et al., 2024) is computed and optimized only at the sampling timesteps within the SDE window $W(l)$.
After each GRPO iteration, the current training steps are checked against the conditions for moving the window, determining whether to shift the window by the window stride $s$.
Algorithm 1 MixGRPO Training Process
1:initial policy model $\pi_{\theta}$; reward models $\{R_{k}\}_{k=1}^{K}$; prompt dataset $\mathcal{C}$; total sampling steps $T$; number of samples per prompt NN;